# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v1

@@ III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
We note that the proof of Proposition V.1 given in Appendix 199 also provides us with closed-form expressions of CRB$(\mathrm{vecs}(\bm{\Sigma}_{0}))$ (187) and $\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0}))$ (197) (199), which are new results.
V-B The Constrained parametric CRB (CCRB) in $\mathcal{P}_{\bm{\nu},g_{0}}$ under $S(\bm{\Sigma}_{0})=s_{0}$
As shown in Appendix 199, the FIM for $\bm{\nu}_{0}\in\Omega$ in the parametric model $\mathcal{P}_{\bm{\nu},g_{0}}$ in (39), i.e. $\mathbf{I}_{\bm{\nu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\nu}_{0}}\mathbf{s}_{\bm{\nu}_{0}}^{T}\}$ in (178) is block-diagonal. Consequently the CCRB on $\bm{\nu}_{0}\in\Omega$ has, in turn, the following block-diagonal structure:
$\mathrm{CCRB}(\bm{\nu}_{0})=\left(\begin{array}[]{cc}\mathrm{CRB}(\bm{\mu}_{0})&\mathbf{0}\\ \mathbf{0}&\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\\ \end{array}\right),\;S(\bm{\Sigma}_{0})=s_{0},$ (100)
where $\mathrm{CRB}(\bm{\mu}_{0})$ is the one already given in (97).
To derive $\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))$, let’s take a step back to the term $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$. In fact, as an alternative method w.r.t. the one used in (98), $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$ can be derived by exploiting the block-structure of the FIM $\mathbf{I}_{\bm{\eta}_{0}}$ in (65). Specifically, from (65)-(71) and by recalling the definition of the efficient FIM $\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$ in (89), we have:
$\begin{split}\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})&\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}-I^{-1}_{s_{0}}\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\mathbf{I}^{T}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\\ &=2^{-1}\alpha(g_{0})\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\frac{(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})}{2}-\frac{1}{m}\mathrm{vec}(\mathbf{I}_{m})\mathrm{vec}(\mathbf{I}_{m})^{T}\right]\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\right]^{T},\\ &=2^{-1}\alpha(g_{0})\mathbf{K}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{K}_{\mathbf{V}_{S,0}},\\ &=2^{-1}\alpha(g_{0})\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{U}_{\mathbf{V}_{S,0}}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1},\end{split}$ (101)
where:
@@
    in the second identity, we used (55) and the two properties of the commutation matrix $\mathbf{K}_{m}$, i.e. i) $\mathbf{K}_{m}(\mathbf{A}\otimes\mathbf{B})=(\mathbf{B}\otimes\mathbf{A})\mathbf{K}_{m}$ and ii) $\mathbf{K}_{m}\mathbf{D}_{m}=\mathbf{D}_{m}$ [23, Ch. 3],
@@
    in the third identity, we introduced the matrix:
    $\mathbf{T}_{\mathbf{A}}\triangleq\mathbf{A}^{-1}\otimes\mathbf{A}^{-1}-m^{-1}\mathrm{vec}(\mathbf{A}^{-1})\mathrm{vec}(\mathbf{A}^{-1})^{T}$ (102)
    for any invertible $m\times m$ matrix $\mathbf{A}$
  3. 3.
    finally, in the forth identity, we used (165) defined in Appendix .6.
Consequently, from the basic properties of the CRB (see e.g. [17, Th. 6.6]) and from (164) in Appendix .6, we have that the CRB on $\mathrm{vecs}(\mathbf{V}_{S,0})$ can be obtained as:
$\begin{split}\mathrm{CRB}(\mathrm{vecs}(\mathbf{V}_{S,0})|s_{0})&=\mathbf{K}_{\mathbf{V}_{S,0}}\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}\mathbf{K}_{\mathbf{V}_{S,0}}^{T}\\ &=\mathbf{U}_{\mathbf{V}_{S,0}}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1}\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\\ &=2\alpha(g_{0})^{-1}\mathbf{U}_{\mathbf{V}_{S,0}}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1}\left[\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{U}_{\mathbf{V}_{S,0}}\right]^{-1}\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1}\\ &=2\alpha(g_{0})^{-1}\mathbf{U}_{\mathbf{V}_{S,0}}\left[\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{U}_{\mathbf{V}_{S,0}}\right]^{-1}\mathbf{U}_{\mathbf{V}_{S,0}}^{T},\end{split}$ (103)
where we used (167) and the fact that $\mathbf{S}_{\mathbf{V}_{S,0}}$ is, by definition, invertible.
Let us now consider the constrained scatter matrix $\bm{\Sigma}_{0}=s_{0}\mathbf{V}_{S,0}$, such that $S(\bm{\Sigma}_{0})=s_{0}$. Using again the properties of the CRB [17, Theo. 6.6] and the fact that, as discussed in the property P5 in Appendix .6, $\mathbf{S}_{\mathbf{V}_{S,0}}=\mathbf{S}_{\bm{\Sigma}_{0}}$ and $\mathbf{U}_{\mathbf{V}_{S,0}}=\mathbf{U}_{\bm{\Sigma}_{0}}$ we have that the CCRB on the $\bm{\Sigma}_{0}$, under the constraint $S(\bm{\Sigma}_{0})=s_{0}$, is given by:
$\begin{split}\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))&=\left[s_{0}\mathbf{K}_{\mathbf{V}_{S,0}}\right]\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}\left[s_{0}\mathbf{K}_{\mathbf{V}_{S,0}}\right]^{T}\\ &=2\alpha(g_{0})^{-1}\mathbf{U}_{\bm{\Sigma}_{0}}\left[\mathbf{U}_{\bm{\Sigma}_{0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\bm{\Sigma}_{0}}\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}_{0}}\right]^{-1}\mathbf{U}_{\bm{\Sigma}_{0}}^{T},\quad\text{for}\;S(\bm{\Sigma}_{0})=s_{0}.\end{split}$ (104)
If we compare (after carefully harmonising the notations) the $\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))$ for $S(\bm{\Sigma}_{0})=s_{0}$ obtained in (104) and the Constrained Semiparametric CRB on $\mathrm{vecs}(\bm{\Sigma}_{0})$ obtained in [22, eq. (59)] we can recognize that they are the same. This surprising result is a direct consequence of the intrinsic geometry of the semiparametric models $\mathcal{P}_{\bm{\eta},g}$ and $\mathcal{P}_{\bm{\nu},g}$ that has been brought to light by Propositions IV.1 and IV.2. To summarize and emphasize the results obtained in this section, let us conclude with the counterpart for the information bounds of Propositions IV.1 and IV.2.
V-C Three equality chains for parametric and semiparametric information bounds in RES distributions
We are finally ready to introduce three sequences of equalities among information bounds for the parameters of interest in RES distributions. The following proposition summarizes the key points from the previous sections and can serve as a take-away message for readers who are not interested in full mathematical details.
Proposition V.2.
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ be a RES distributed vector with location vector $\bm{\mu}_{0}$, scatter matrix $\bm{\Sigma}_{0}$ and density generator $g_{0}\in\bar{\mathcal{G}}$. Let $S(\bm{\Sigma}_{0})=s_{0}$ be a constraint on the covariance matrix where $S$ is a matrix function satisfying Assumptions A1, A2 and A3 and let $\mathbf{V}_{S,0}=s_{0}^{-1}\bm{\Sigma}_{0}$ the related shape matrix. Moreover:
  • Let $\mathbf{D}_{S}^{\bm{\Sigma}_{0}}\triangleq\frac{\partial S(\bm{\Sigma}_{0})}{\partial\bm{\Sigma}_{0}}$,
  • Let $\mathbf{U}_{\bm{\Sigma}_{0}}$ be a matrix whose columns form an orthonormal basis for the null space of $\mathrm{vecs}(\mathbf{D}_{S}^{\bm{\Sigma}_{0}})^{T}$, i.e.:
    $\mathrm{vecs}(\mathbf{D}_{S}^{\bm{\Sigma}_{0}})^{T}\mathbf{U}_{\bm{\Sigma}_{0}}=\mathbf{0},\quad\mathbf{U}_{\bm{\Sigma}_{0}}^{T}\mathbf{U}_{\bm{\Sigma}_{0}}=\mathbf{I}_{m(m+1)/2-1}.$
  • $\mathbf{T}_{\bm{\Sigma}_{0}}\triangleq\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1}-m^{-1}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}$,
Then, the following chain of equalities hold:
  1. 1.
    Relation among the SCRB for the location vector $\bm{\mu}_{0}$ in the semiparametric model $\mathcal{P}_{\bm{\nu},g}$ (or $\mathcal{P}_{\bm{\eta},g}$) and three different CRB in the parametric models $\mathcal{P}_{\bm{\nu},s}$ (or $\mathcal{P}_{\bm{\eta},s}$), $\mathcal{P}_{\bm{\nu},s,\bm{\zeta}}$ (or $\mathcal{P}_{\bm{\eta},s,\bm{\zeta}}$) and $\mathcal{P}_{\bm{\nu}}$ (or $\mathcal{P}_{\bm{\eta}}$):
    $\begin{split}&\beta(g_{0})^{-1}\bm{\Sigma}_{0}\\ &=\mathrm{SCRB}(\bm{\mu}_{0}|s_{0},g_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally unknown}\right]\\ &=\mathrm{CRB}(\bm{\mu}_{0}|s_{0},\bm{\zeta}_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally known up to its parameters}\;\bm{\zeta}_{0}\right]\\ &=\mathrm{CRB}(\bm{\mu}_{0}|s_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{fully known}\right]\\ &=\mathrm{CRB}(\bm{\mu}_{0}),\quad\left[s_{0}\;\text{known},\;g_{0}\;\text{fully known}\right].\end{split}$ (105)
  2. 2.
    Relation among the constrained SCRB (CSCRB) for the constrained scatter matrix $\bm{\Sigma}_{0}$ in the semiparametric model $\mathcal{P}_{\bm{\nu},g}$ and three different constrained CRB (CCRB) in the parametric models $\mathcal{P}_{\bm{\nu},s}$, $\mathcal{P}_{\bm{\nu},s,\bm{\zeta}}$ and $\mathcal{P}_{\bm{\nu}}$, under $S(\bm{\Sigma}_{0})=s_{0}$:
    $\begin{split}2\alpha(g_{0})^{-1}&\mathbf{U}_{\bm{\Sigma}_{0}}\left[\mathbf{U}_{\bm{\Sigma}_{0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\bm{\Sigma}_{0}}\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}_{0}}\right]^{-1}\mathbf{U}_{\bm{\Sigma}_{0}}^{T},\quad\text{for}\;S(\bm{\Sigma}_{0})=s_{0}.\\ &=\mathrm{CSCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})|s_{0},g_{0}),\;\forall S(\bm{\Sigma}_{0})=s_{0}\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionallyunknown}\right]\\ &=\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})|s_{0},\bm{\zeta}_{0}),\;\forall S(\bm{\Sigma}_{0})=s_{0}\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally known up to its parameters}\;\bm{\zeta}_{0}\right]\\ &=\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})|s_{0}),\;\forall S(\bm{\Sigma}_{0})=s_{0}\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{fully known}\right]\\ &=\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})),\;\textit{iff}\;|\bm{\Sigma}_{0}|^{1/m}=s_{0}\quad\left[s_{0}\;\text{known},\;g_{0}\;\text{fully known}\right].\end{split}$ (106)
  3. 3.
@@ V-C Three equality chains for parametric and semiparametric information bounds in RES distributions
    $\begin{split}\alpha^{-1}(g_{0})&\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},\\ &=\mathrm{SCRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0},g_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally unknown}\right]\\ &=\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0},\bm{\zeta}_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally known up to its parameters}\;\bm{\zeta}_{0}\right]\\ &=\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{fully known}\right]\\ &=\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0})),\;\textit{iff}\;|\mathbf{V}_{S_{d},0}|^{1/m}=1\quad\left[s_{0}\;\text{known},\;g_{0}\;\text{fully known}\right].\end{split}$ (107)
@@
The proof is concluded by noticing that the SCRB and the different CRBs are defined as the inverse of the related FIM according to the Hájek-Le Cam convolution theorem for parametric models [1, Sect. 2.3, Th. 1], semiparametric models [1, Sect. 3.3, Theo. 2] and their constrained counterparts [27]. ∎
@@ VI-A Some preliminaries
Let us start with some preliminaries. Let $\mathcal{X}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g)$ be a RES-distributed random vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})\in\mathbb{R}^{m}$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})\in\mathcal{S}_{m}^{\mathbb{R}}$ are parameterized by a $d$-dimensional parameter vector $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Theta\triangleq\Gamma\times\Psi$. As in Sect. II, $\Gamma\subseteq\mathbb{R}^{q}$ denotes the set of the (finite-dimensional) parameter vectors $\bm{\gamma}$ of interests, $\Psi\subseteq\mathbb{R}^{r}$ denotes the set of (finite-dimensional) nuisance parameter vectors $\bm{\xi}$ and $\bar{\mathcal{G}}$ is the set of the (infinite-dimensional) nuisance functions $g$ already defined in (34). In the following, we always assume that the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ satisfies the following assumptions:
@@
    the inverse $[\bm{\Sigma}(\bm{\theta})]^{-1}$ exists for all $\bm{\theta}\in\Theta$.
@@
Let us start with the calculation of the score function $\mathbf{s}_{\bm{\theta}_{0}}$. Following the derivation in [28, Sec. 3.1] and [29, Sec. III], each entry of $\mathbf{s}_{\bm{\theta}_{0}}$ can be easily evaluated as:
@@
$[\mathbf{J}[\bm{\mu}_{0}]]_{i,j}=\left.\frac{\partial[\bm{\mu}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}},\;[\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]]_{i,j}=\left.\frac{\partial[\mathrm{vec}(\bm{\Sigma}(\bm{\theta}))]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$ (120)

### v2

@@ V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\eta},g}$
[…]
V-B Equality chains for parametric and semiparametric information bounds in RES distributions
We are finally ready to introduce two sequences of equalities among information bounds for the parameters of interest in RES distributions. The following proposition summarizes the key points from the previous sections and can serve as a take-away message for readers who are not interested in full mathematical details.
Proposition V.2.
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ be a RES distributed vector with location vector $\bm{\mu}_{0}$, scatter matrix $\bm{\Sigma}_{0}$ and density generator $g_{0}\in\bar{\mathcal{G}}$. Let $S(\bm{\Sigma}_{0})=s_{0}$ be a constraint on the covariance matrix where $S$ is a matrix function satisfying Assumptions A1, A2 and A3 and let $\mathbf{V}_{S,0}=s_{0}^{-1}\bm{\Sigma}_{0}$ the related shape matrix. Then, the following chain of equalities hold:
@@ V-B Equality chains for parametric and semiparametric information bounds in RES distributions
    Relation among the SCRB for the location vector $\bm{\mu}_{0}$ in the semiparametric model $\mathcal{P}_{\bm{\eta},g}$ and three different CRB in the parametric models $\mathcal{P}_{\bm{\eta},s}$, $\mathcal{P}_{\bm{\eta},s,\bm{\zeta}}$ and $\mathcal{P}_{\bm{\eta}}$:
[…]
@@
[…]
    where $\mathbf{P}_{S}(\mathbf{V}_{S,0})\triangleq\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0})\bm{\nabla}_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})$.
@@
The proof is concluded by noticing that the SCRB and the different CRBs are defined as the inverse of the related FIM according to the Hájek-Le Cam convolution theorem for parametric models [1, Sect. 2.3, Th. 1], semiparametric models [1, Sect. 3.3, Theo. 2]. ∎
@@ VI-A Some preliminaries
Let $\mathcal{X}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g)$ be a RES-distributed random vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})\in\mathbb{R}^{m}$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})\in\mathcal{S}_{m}^{\mathbb{R}}$ are parameterized by a $d$-dimensional parameter vector $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Theta\triangleq\Gamma\times\Psi$. As in Sect. II, $\Gamma\subseteq\mathbb{R}^{q}$ denotes the set of the (finite-dimensional) parameter vectors $\bm{\gamma}$ of interests, $\Psi\subseteq\mathbb{R}^{r}$ denotes the set of (finite-dimensional) nuisance parameter vectors $\bm{\xi}$ and $\bar{\mathcal{G}}$ is the set of the (infinite-dimensional) nuisance functions $g$ already defined in (34). In the following, we always assume that the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ satisfies the following assumptions:
@@
    the inverse $\bm{\Sigma}(\bm{\theta})]^{-1}$ exists for all $\bm{\theta}\in\Theta$.
@@
Let us start with the calculation of the score function $\mathbf{s}_{\bm{\theta}_{0}}$. Following the derivation in [26, Sec. 3.1] and [27, Sec. III], each entry of $\mathbf{s}_{\bm{\theta}_{0}}$ can be easily evaluated as:
@@
$\mathbf{J}[\bm{\mu}_{0}]]_{i,j}=\left.\frac{\partial[\bm{\mu}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}},\;[\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]]_{i,j}=\left.\frac{\partial[\mathrm{vec}(\bm{\Sigma}(\bm{\theta}))]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$ (114)